The area of a sector of a circle is 44cm. What is the radius of the circle if the angle at the centre of the circle is 140?
Answer by nerdybill(7384)

The area of a sector of a circle is 44cm. What is the radius of the circle if the angle at the centre of the circle is 140?
.
sectorArea = n/360 * (pi)r^2
where
sectorArea is 44
n is number of degrees (140)
pi is 3.14
r is radius
.
44 = 140/360 * (3.14)r^2
44 = 7/18 * (3.14)r^2
44(18/7) = (3.14)r^2
44(18/7) = (3.14)r^2
113.1429 = (3.14)r^2
36.0328 = r^2
6 cm = r